# Multi-Constraint Safe RL with Objective Suppression for Safety-Critical Applications

### v2

@@ Abstract
Safe reinforcement learning tasks with multiple constraints are a challenging domain despite being very common in the real world. In safety-critical domains, properly handling the constraints becomes even more important. To address this challenge, we first describe the multi-constraint problem with a stronger Uniformly Constrained MDP (UCMDP) model; we then propose Objective Suppression, a novel method that adaptively suppresses the task reward maximizing objectives according to a safety critic, as a solution to the Lagrangian dual of a UCMDP. We benchmark Objective Suppression in two multi-constraint safety domains, including an autonomous driving domain where any incorrect behavior can lead to disastrous consequences. Empirically, we demonstrate that our proposed method, when combined with existing safe RL algorithms, can match the task reward achieved by our baselines with significantly fewer constraint violations.
@@ 1 Introduction
Reinforcement learning (RL) is a general approach to solving challenging tasks in many domains such as robotics, navigation, and even generative modeling. Policies learned with RL seek to maximize a “task reward” over a long run which can be specified either via manually defined functions or through learned models. However, without careful tuning of this reward function, it can be difficult for policies to learn to perform well in safety-critical situations like those in the autonomous driving domain. In order to reliably use RL policies in safety-critical situations, we consider the use of constrained RL in order to prevent the policy from exhibiting dangerous behavior. Most prior works on constrained or safe RL tend to put an emphasis on single-constraint violations; however, in most real-world settings there are often multiple constraints that can even be conflicting, which is under-explored in prior work. Take for instance the case of autonomous driving, two very simple constraints are (1) avoiding collision and (2) maintaining a buffer distance from static objects. In some cases, it is impossible to satisfy both constraints while also simultaneously making progress along the ego route.
Existing safe RL works rarely focus on the multi-constraint issue. Methods that rely on linearly combining the task reward and constraints can struggle to assign a set of weights for all the constraints without some of the constraints being overshadowed by others; hierarchical methods, on the other hand, can face difficulties in building multiple hierarchies. On the other hand, safety-critical domains pose extra demands on the behavioral safety. Previously, safe RL algorithms typically model the constrained RL problem as a constrained Markov Decision Process (CMDP, Altman (1999); García & Fernández (2015); Achiam et al. (2017)), which only constrains the behaviors in expectation. With the CMDP constraints, less frequently visited states are prone to having higher constraint violation probabilities, which can be dangerous in safety-critical scenarios. In this work, we instead propose to use a stronger uniformly constrained Markov Decision Process (UCMDP), where we put constraints uniformly on the support of visiting distribution of state and actions.
In light of the challenges, we present Objective Suppression, a safe RL algorithm derived from the reward-switching objectives that makes adaptive choices of suppressing and balancing the task reward maximizing objective and constraint-satisfying objectives, as a solution to the UCMDP model. Objective Suppression can also be easily combined with existing safe RL approaches as a compound constraint-enforcing method. As an example, combining with hierarchical methods like Recovery RL would improve constraint satisfaction property during training instead of only in convergence. Moreover, combining multiple perpendicular methods is also helpful to enforce multiple constraints.
@@
We empirically test our method in two challenging domains featuring multiple constraints: a Mujoco-Ant (Todorov et al., 2012; Brockman et al., 2016) domain with dynamic obstacles, where our method lowers the number of collisions by 33%; and the Safe Bench (Xu et al., 2022) domain, where our method reduces the constraint violations by at least half. In both domains, our method achieves the results without significant sacrifice of task reward.
@@ 2 Related Works
Lagrangian Methods
Lagrangian relaxation (Altman, 1999; Bertsekas, 2016) uses a primal-dual method to turn the constrained optimization problem of safe RL into an unconstrained one. Chow et al. (2018) demonstrates Lagrangian relaxation can be adopted to safe RL problems and achieve satisfying empirical performances. RCPO (Tessler et al., 2019) turns Lagrangian multipliers into reward penalty weights to reach constrained goals with both theoretical and empirical evidence. Stooke et al. (2020) takes Lagrangian methods into controls domain by utilizing the derivatives of the constraint function.
Hierarchical Methods
Another way of solving safe RL problems is to apply a safety layer (Dalal et al., 2018) on top of the task reward-maximizing policy. Dalal et al. (2018) derives a closed-form solution for action correction by learning a linearized model. Recovery RL (Thananjeyan et al., 2020) learns a parameterized recovery policy that leads the agent away from dangerous areas. Bharadhwaj et al. (2021) takes inspiration from offline RL algorithms and proposes a hierarchical method with a dynamic error tolerance threshold that achieves provable per-iteration safety guarantees.
@@
Most prior work in learning-based methods in AV relies on behavior cloning approaches (Hu et al., 2023; Shao et al., 2023). Some methods add layers of hierarchy (Bronstein et al., 2022) in order to learn improved BC policies but can still struggle with safety. Pure cloning-based approaches have been demonstrated to be insufficient to handle long-tail cases or ensure safety (Lu et al., 2022; Gulino et al., 2023) which motivates the exploration of RL-based approaches.
@@ 3 Preliminaries
We describe a safe RL problem under the assumption of a constrained Markov Decision Process (CMDP, Altman (1999); García & Fernández (2015); Achiam et al. (2017)), where the agent is required to maximize its expected return while ensuring all the safety constraints are met. Formally, a CMDP is defined as a 7-tuple $(\mathcal{S},\mathcal{A},P,R,\gamma,\mathcal{C},\epsilon,\mu_{0})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the stochastic state transition function, $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the task reward function, $\gamma$ is the discount factor, $\mathcal{C}=\{(C_{i}:\mathcal{S}\rightarrow\{0,1\},\gamma_{C_{i}})\}$ is the set of constraints and their corresponding discount factors, ϵ+italic-$\epsilon\in\mathbb{R}^{+}$ is the constraint violation limit, and $\mu_{0}:\mathcal{S}\rightarrow[0,1]$ is the initial state distribution. Let $n=|\mathcal{C}|$ denote the number of constraints. In the context of safe RL, these constraints are also risks.
@@
One regime of solving CMDPs is to apply a safety layer to adjust the unsafe actions Dalal et al. (2018); Thananjeyan et al. (2020); Yu et al. (2022). The safety layer overwrites the proposed actions that are deemed unsafe according to the safety critic $Q_{C_{i}}^{\pi}$:
@@
$\sigma:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is a safety-ensuring policy. One choice of $\sigma$ is to project the actions into a feasible space (Pham et al., 2017; Dalal et al., 2018), e.g., $\sigma(s)=\arg\min_{a}\|a-\tilde{a}\|\ \text{s.t.}\ Q_{C_{i}}^{\pi}(s_{t},a)\leq\epsilon\ \text{for all}\ i\in\{1,\dots,n\}$, where $\tilde{a}\sim\pi(s)$. Another choice is to use a separate parameterized policy (Thananjeyan et al., 2020; Yu et al., 2022). Recovery RL (Thananjeyan et al., 2020) trains a recovery policy to minimize the constraint violations, i.e., $\sigma^{*}=\arg\min_{\sigma}\sum_{i=1}^{n}w_{i}\mathcal{J}_{C_{i}}^{(\pi,\sigma)}$.
@@
In this work we consider parameterized policies denoted as $\pi_{\theta}$. The derived gradients (Williams, 1992) of $\mathcal{J}_{R}^{\pi_{\theta}}$ and $\mathcal{J}_{C_{i}}^{\pi_{\theta}}$ are:
@@ 4 Objective Suppression
We observe that previous safe RL algorithms struggle in multi-constraint scenarios. We propose a new method to enforce safety constraints on policy optimization based on adaptively suppressing the task reward objectives that combines well with other safe RL algorithms such as Recovery RL in multiple constraint scenarios. We first formulate our safety-critical problem setting with a stronger model than CMDPs, the uniformly constrained MDPs (UCMDPs) in Section 4.1. In Section 4.2, we motivate our proposed Objective Suppression method from a reward-switching perspective, as a non-parametric approximation to the solution of the Lagrangian dual of UCMDPs. Finally, in Section 4.3, we discuss the potential of combining Objective Suppression with other safe RL algorithms.
@@ 4.1 Uniformly Constrained MDPs
CMDPs constraint behaviors in expectation. However, in safety-critical domains such as autonomous driving, safety guarantees in less frequently visited situations should be treated with equal attention. In light of this, we propose a stronger version of CMDPs, named uniformly constrained MDPs (UCMDPs). Uniformly constrained MDPs share the same definition as the CMDPs except for the optimization objective constraint, where we pose a uniform one instead of the expectation one in CMDPs:
@@
$d^{\pi}(s,a)$ is the distribution of state-action pairs under ${\pi}$. The Lagrangian dual problem of (14) is $\min_{\lambda\geq 0}\max_{\theta}\mathcal{L}(\theta,\lambda)$, where
@@
which we further relax to
@@ 4.2 Objective Suppression
To enforce safety constraints on policy optimization, we would like to train a policy that automatically switches between optimizing for task reward objective $\mathcal{J}_{R}^{\pi}$ and risk minimization objective $\mathcal{J}_{C_{i}}^{\pi}$. One way to accomplish this is to switch the optimization objective in hindsight (Zhou et al., 2022). Specifically, at step $t$ of a trajectory $\tau$, if any risk is encountered after $t$, we then switch to the risk minimization objective; otherwise, we remain using the original task reward objective. Let $z_{t}^{i}(\tau)={\mathbbm{1}}_{\{\text{$\tau$ encounters risk $i$ after step $t$}\}}$ and $z_{t}^{-}(\tau)={\mathbbm{1}}_{\{\text{$\tau$ does not encounter any risk after step $t$}\}}=\prod_{i}1-z_{t}^{i}(\tau)$. We define the hard-switching objective as
@@
where $w_{1},\dots,w_{n}$ are the weights to balance the different risk minimization objectives. The hard-switching objective (22) is an interpolation of $\nabla_{\theta}\mathcal{J}_{R}^{\pi}$ and $\nabla_{\theta}\mathcal{J}_{C_{i}}^{\pi}$. However, this objective faces problems in practice. For one, the hard-switching between multiple objectives raises the variance of the gradient, which is widely acknowledged to have a negative impact on training; for another, balancing the different objectives with weights introduces a new set of hyperparameters. If handled improperly, some of the objectives will become dominant over others and can result in disastrous outcomes.
To deal with the aforementioned two problems, we rewrite (22) into:
@@ 5.1 Environments and baselines
This environment is adapted from the 8-DOF ant-v4 environment from gym (Brockman et al., 2016) Mujoco (Todorov et al., 2012). The main task of the agent is to control the ant to reach a target point. There are two constraints in this environment. The first constraint is to avoid randomly spawned obstacles and the second is to avoid collapsing.
@@
This environment is a CARLA (Dosovitskiy et al., 2017) based benchmark for safety in various different driving scenarios (Xu et al., 2022). The task reward comes from making progress along the designated route while the constraints come from (1) collisions with obstacles (e.g., pedestrians, curbs, etc) which will terminate the episode, and (2) leaving the lane, i.e. lateral deviation from the lane of travel. The policy is given access to a bird-eve-view rendering of the scene in addition to a 4-D observation space covering lane placement, speed, and the distances to objects. The action space is a continuous 2-D space consisting of control parameters: acceleration and steering angle.
@@
We test our objective suppression method in the two aforementioned environments, Safe Mujoco-Ant and Safe Bench. We also implement two baselines, a naive Reward Penalty baseline, where the optimization objective is a fine-tuned weighted sum of the task reward objective and risk minimization objective; and Recovery RL (Thananjeyan et al., 2020).

### v3

# Uniformly Safe RL with Objective Suppression for Multi-Constraint Safety-Critical Applications

@@ Abstract
Safe reinforcement learning tasks are a challenging domain despite being very common in the real world. The widely adopted CMDP model constrains the risks in expectation, which makes room for dangerous behaviors in long-tail states. In safety-critical domains, such behaviors could lead to disastrous outcomes. To address this issue, we first describe the problem with a stronger Uniformly Constrained MDP (UCMDP) model where we impose constraints on all reachable states; we then propose Objective Suppression, a novel method that adaptively suppresses the task reward maximizing objectives according to a safety critic, as a solution to the Lagrangian dual of a UCMDP. We benchmark Objective Suppression in two multi-constraint safety domains, including an autonomous driving domain where any incorrect behavior can lead to disastrous consequences. On the driving domain, we evaluate on open source and proprietary data and evaluate transfer to a real autonomous fleet. Empirically, we demonstrate that our proposed method, when combined with existing safe RL algorithms, can match the task reward achieved by baselines with significantly fewer constraint violations.
Keywords: Safe Reinforce Learning, Autonomous Vehicles
@@ 1 Introduction
Reinforcement learning (RL) is a general approach to solving challenging tasks in many domains such as robotics, navigation, and even generative modeling [1, 2, 3, 4]. Policies learned with RL seek to maximize a “task reward” over a long run which can be specified either via manually defined functions or through learned models. However, without careful tuning of this reward function, it can be difficult for policies to learn to perform well in safety-critical situations like those in the autonomous driving domain [5, 4]. Safe RL [6] attempts to solve this by modeling RL with safety concerns as a constrained optimization problem and has been widely adopted in safety-critical RL problems [7, 8]
One common practice of safe RL is to use the Constrained Markov Decision Process (CMDP, [9, 10]) model. CMDPs characterize safety concerns as state-action-dependent scalar constraint functions, similar to the task reward. The safety guarantee is then realized by constraining the expected return of the constraint functions expressed as [8]:
$C_{i}:J_{C_{i}}(\pi)=\mathbb{E}_{\tau\sim\pi}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{\infty}\gamma^{t}C_{i}(s_{t},a_{t},s_{t+1})}\right]$ (1)
However, one drawback of this formulation is that constraining only the expected return might lead to high-risk behaviors in less frequently visited states according to $\pi$, which can lead to disastrous outcomes in real-world safety-critical domains. Addressing this issue is particularly important in domains such as self-driving cars, where it is crucial not to compromise critical safety constraints (e.g., prioritizing human safety above all) for the sake of achieving task-level objectives, even in rare long-tail catastrophic events.
We thus propose Uniformly Constrained MDP (UCMDP), a variant of CMDP which removes the expectation imposed by the task policy $\pi$ in the constraints objective in Equation 1 by imposing constraints uniformly across all reachable states. This reformulation effectively tightens the safety constraints to ensure optimization toward worst-case scenarios and rare long-tail events. To solve the UCMDP, we use primal-dual methods [11], transforming UCMDP into its Lagrangian dual form, with state-dependent Lagrangian multipliers.
We then derive a non-parametric approach that approximates the state-dependent Lagrangian multipliers from objective-switching methods [12]. The resulting algorithm is our proposed Objective Suppression, a safe RL algorithm that makes adaptive choices of suppressing and balancing the task reward maximizing objective and constraint-satisfying objectives, as a solution to the UCMDP model. Furthermore, the primal-dual nature of Objective Suppression allows us to organically combine it with existing hierarchical safe-RL methods such as Recovery RL [13].
In multi-constraint safe RL domains, methods that rely on linearly combining the task reward and constraints can struggle to assign a set of weights for all the constraints without some of the constraints being overshadowed by others; hierarchical methods, on the other hand, can face difficulties in building multiple hierarchies. We then proceed to show that by combining Objective Suppression with hierarchical methods, we can utilize both regimes to accommodate different sources of safety constraint.
@@
We empirically test our method in two challenging domains featuring multiple constraints: a Mujoco-Ant [14, 15] domain with dynamic obstacles, where our method lowers the number of collisions by 33%; and the Safe Bench [16] domain, where our method reduces the constraint violations by at least half. In both domains, our method achieves the results without significant sacrifice of task return. We also discuss our experiments conducted on both open-source and proprietary data, and evaluate their transfer to a real autonomous fleet.
We summarize our core contributions as follows:
  • We propose UCMDP, a variant of CMDP that puts uniform attention on all states to prevent dangerous behaviors in long-tail cases.
  • We propose Objective Suppresion, a practical algorithm that solves UCMDPs that can be applied to existing safe RL methods.
  • We adapt two multi-constraint safe RL benchmarks with continuous state and action space from Mujoco-Ant and Safe Bench. We then empirically show that Objective Suppresion significantly improves safety in the proposed benchmarks.
@@ 2 Related Works
Lagrangian Methods for CMDPs
Lagrangian relaxation [9, 17] uses a primal-dual method to turn the constrained optimization problem of safe RL into an unconstrained one. Chow et al. [11] demonstrates Lagrangian relaxation can be adopted to safe RL problems and achieve satisfying empirical performances. RCPO [18] turns Lagrangian multipliers into reward penalty weights to reach constrained goals with both theoretical and empirical evidence. Stooke et al. [19] takes Lagrangian methods into controls domain by utilizing the derivatives of the constraint function. Objective Suppression solves an approximation of the Lagrangian dual of UCMDP.
Hierarchical Methods for CMDPs
Another way of solving safe RL problems is to apply a safety layer [20] on top of the task reward-maximizing policy. Dalal et al. [20] derives a closed-form solution for action correction by learning a linearized model. Recovery RL [13] learns a parameterized recovery policy that leads the agent away from dangerous areas. Bharadhwaj et al. [21] takes inspiration from offline RL algorithms and proposes a hierarchical method with a dynamic error tolerance threshold that achieves provable per-iteration safety guarantees.
@@
Most prior work in learning-based methods in AV relies on behavior cloning approaches [22, 23]. Some methods add layers of hierarchy [24] in order to learn improved BC policies but can still struggle with safety. Pure cloning-based approaches have been demonstrated to be insufficient to handle long-tail cases or ensure safety [25, 26] which motivates the exploration of RL-based approaches.
@@ 3 Preliminaries
We describe a safe RL problem under the assumption of a constrained Markov Decision Process (CMDP, [9, 27, 10]), where the agent is required to maximize its expected return while ensuring all the safety constraints are met. Formally, a CMDP is defined as a 7-tuple $(\mathcal{S},\mathcal{A},P,R,\gamma,\mathcal{C},\epsilon,\mu_{0})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the stochastic state transition function, $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the task reward function, $\gamma$ is the discount factor, $\mathcal{C}=\{(C_{i}:\mathcal{S}\rightarrow\{0,1\},\gamma_{C_{i}})\}$ is the set of constraints and their corresponding discount factors, ϵ+italic-$\epsilon\in\mathbb{R}^{+}$ is the constraint violation limit, and $\mu_{0}:\mathcal{S}\rightarrow[0,1]$ is the initial state distribution. Let $n=|\mathcal{C}|$ denote the number of constraints. In the context of safe RL, these constraints are also risks.
@@
One regime of solving CMDPs is to apply a safety layer to adjust the unsafe actions [20, 13, 28]. The safety layer overwrites the proposed actions that are deemed unsafe according to the safety critic $Q_{C_{i}}^{\pi}$:
@@
$\sigma:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is a safety-ensuring policy. One choice of $\sigma$ is to project the actions into a feasible space [29, 20], e.g., $\sigma(s)=\arg\min_{a}\|a-\tilde{a}\|\ \text{s.t.}\ Q_{C_{i}}^{\pi}(s_{t},a)\leq\epsilon\ \text{for all}\ i\in\{1,\dots,n\}$, where $\tilde{a}\sim\pi(s)$. Another choice is to use a separate parameterized policy [13, 28]. Recovery RL [13] trains a recovery policy to minimize the constraint violations, i.e., $\sigma^{*}=\arg\min_{\sigma}\sum_{i=1}^{n}w_{i}\mathcal{J}_{C_{i}}^{(\pi,\sigma)}$.
@@
In this work we consider parameterized policies denoted as $\pi_{\theta}$. The derived gradients [30] of $\mathcal{J}_{R}^{\pi_{\theta}}$ and $\mathcal{J}_{C_{i}}^{\pi_{\theta}}$ are:
@@ 4 Objective Suppression
We propose a new method to enforce safety constraints on policy optimization based on adaptively suppressing the task reward objectives that combines well with other safe RL algorithms such as Recovery RL in multiple constraint scenarios. We first formulate our safety-critical problem setting with a stronger model than CMDPs, the Uniformly Constrained MDPs (UCMDPs) in Section 4.1. In Section 4.2, we motivate our proposed Objective Suppression method from a reward-switching perspective, as a non-parametric approximation to the solution of the Lagrangian dual of UCMDPs. In Section 4.3, we discuss the potential of combining Objective Suppression with other safe RL algorithms. Finally, in Section 4.4, we make a connection from Objective Suppression to Hierarchical RL to provide more insights to our method.
@@ 4.1 Uniformly Constrained MDPs
CMDPs constraint behaviors in expectation. However, in safety-critical domains such as autonomous driving, safety guarantees in less frequently visited situations should be treated with equal attention. In light of this, we propose a stronger version of CMDPs, named Uniformly Constrained MDPs (UCMDPs). UCMDPs share the same definition as the CMDPs except for the optimization objective constraint, where we pose a uniform one instead of the expectation one in CMDPs:
@@
$d^{\pi}(s,a)$ is the distribution of state-action pairs under ${\pi}$.
Comparison with SCMDPs
A similar model that focuses on state-wise constraints is State-wise Constrained MDP (SCMDP, [31]). UCMDPs differ from SCMDPs in that UCMDPs constrain the policy-dependent expected return of constraint violations ($Q_{C_{i}}^{\pi}$) instead of a purely state-based constraint function. In dynamic environments such as driving, there are often states that are inherently risky due to other actors or environment conditions. As an example, consider a narrow corridor where two vehicles need to negotiate. Entering the corridor is inherently more risky than yielding due to the uncertainty over the behavior of the other agent. In the SCMDP formulation, we would disallow entering the region with a pure state-based constraint function resulting in over conservative exploration. However, in the UCMDP formulation, the policy would be allowed to explore the region and discover a safe negotiation of the narrow corridor.
Lagrangian dual of UCMDPs
The Lagrangian dual problem of (15) is
$\min_{\lambda\geq 0}\max_{\theta}\mathcal{L}(\theta,\lambda),$ (16)
where
@@
which we further relax to:
@@ 4.2 Objective Suppression
To enforce safety constraints on policy optimization, we would like to train a policy that automatically switches between optimizing for task reward objective $\mathcal{J}_{R}^{\pi}$ and risk minimization objective $\mathcal{J}_{C_{i}}^{\pi}$. One way to accomplish this is to switch the optimization objective in hindsight [12]. Specifically, at step $t$ of a trajectory $\tau$, if any risk is encountered after $t$, we then switch to the risk minimization objective; otherwise, we remain using the original task reward objective. Let $z_{t}^{i}(\tau)={\mathbbm{1}}_{\{\text{$\tau$ encounters risk $i$ after step $t$}\}}$ and $z_{t}^{-}(\tau)={\mathbbm{1}}_{\{\text{$\tau$ does not encounter any risk after step $t$}\}}=\prod_{i}1-z_{t}^{i}(\tau)$. We define the hard-switching objective as
@@
where $w_{1},\dots,w_{n}$ are the weights to balance the different risk minimization objectives. The hard-switching objective (24) is an interpolation of $\nabla_{\theta}\mathcal{J}_{R}^{\pi}$ and $\nabla_{\theta}\mathcal{J}_{C_{i}}^{\pi}$. However, this objective faces problems in practice. For one, the hard-switching between multiple objectives raises the variance of the gradient, which is widely acknowledged to have a negative impact on training; for another, balancing the different objectives with weights introduces a new set of hyperparameters. If handled improperly, some of the objectives will dominate others and can result in disastrous outcomes. To deal with the aforementioned two problems, we rewrite (24) into:
@@ 4.3 Combining with Existing Safe RL Algorithms
4.4 Connection to the Hierarchical Reinforcement Learning (HRL) Framework
An alternative perspective on our approach is to consider it as a hierarchical reinforcement learning model, specifically within the framework of options [32]. An option $\mathcal{O}\equiv\langle\mathcal{I}_{o},\pi_{o},\beta_{o}\rangle$ defines a temporally extended action (or a policy) where $\mathcal{I}_{o}(s)$ denotes the set of states where the option can be initiated, $\pi_{o}(s,a)\mapsto[0,1]$ denotes the (stochastic) policy associated with that option, and $\beta(s)\mapsto[0,1]$ denotes the termination condition, or the probability of terminating the option in a given state $s$.
In our framework, we can view the high level policy $\mathcal{J}_{switch}^{\pi}$ as a stochastic switching strategy that selects between two options: $\mathcal{O}_{task}\equiv\langle\mathcal{I}_{task},\pi_{R},\beta_{task}\rangle$ and $\mathcal{O}_{recovery}\equiv\langle\mathcal{I}_{recovery},\pi_{C},\beta_{recovery}\rangle$. For option $\mathcal{O}_{task}$, the initiation set $\mathcal{I}_{task}$ denotes the set of risk free states where the agent can safely optimize the task reward, the policy $\pi_{R}$ denotes the task policy, and the termination condition $\beta_{task}$ gets triggered in any state where any of the constraints (risks) are violated. For option $\mathcal{O}_{recovery}$, the initiation set $\mathcal{I}_{recovery}$ denotes the set of states at which at least one constraint is violated, the policy $\pi_{C}$ denotes the recovery policy, and the termination condition $\beta_{recovery}$ gets triggered when no constraints (risks) are violated and the agent can safely execute the task policy until the next safety violation. In this view, our framework essentially learns the model of the options $\mathcal{O}_{task}$ and $\mathcal{I}_{recovery}$ in terms of their policies and termination conditions.
@@ 5.1 Environments and baselines
This environment is adapted from the 8-DOF ant-v4 environment from gym [15] Mujoco [14]. The main task of the agent is to control the ant to reach a target point. There are two constraints in this environment. The first constraint is to avoid randomly spawned obstacles and the second is to avoid collapsing.
@@
This environment is a CARLA [33] based benchmark for safety in various different driving scenarios [16]. The task reward comes from making progress along the designated route while the constraints come from (1) collisions with obstacles (e.g., pedestrians, curbs, etc) which will terminate the episode, and (2) leaving the lane, i.e. lateral deviation from the lane of travel. The policy is given access to a bird-eve-view rendering of the scene in addition to a 4-D observation space covering lane placement, speed, and the distances to objects. The action space is a continuous 2-D space consisting of control parameters: acceleration and steering angle.
@@
We test our objective suppression method in the two aforementioned environments, Safe Mujoco-Ant and Safe Bench. We also implement two baselines, a naive Reward Penalty baseline, where the optimization objective is a fine-tuned weighted sum of the task reward objective and risk minimization objective; and Recovery RL [13].